OpenStudy (anonymous):

\[\int\limits_{-3}^{9}e^{xIn3}dx\]

sam (.sam.):

Let u=xln3 dx=(du)/(ln3) (e^(u))/(ln3) (e^(xln3))/(ln3)+C (e^((9)ln3))/(ln3)-[(e^((-3)ln3))/(ln3)] 531440/27ln3

OpenStudy (anonymous):

The inside if just 3^x even before integration if you simplify it

OpenStudy (anonymous):

The answer is \[{1 \over In3}(3^9-{1\over27}) ~~~~~ In~~this~~form\]

OpenStudy (anonymous):

Thanks Sam

OpenStudy (anonymous):

[(3^x/ln3)-(3^x/ln3)] simplify out 1/ln3 1/ln3 [(3^x)-(3^x)] replace the first x with 9 and the second with -2

OpenStudy (anonymous):

integral of 3^x = 3^x/ln3

OpenStudy (anonymous):

Thanks. I understand :)